# All-Sky Kinematics of the Distant Halo: The Reflex Response to the LMC

### v1

@@ 4.1 Global Velocity Model with Reflex Motion
The direction of $v_{\mathrm{travel}}$ can be encoded by any on-sky angular quantity. Although PP21 used the Galactic coordinates $\ell{}$ and $b$ directly, we find that this leads to inefficient convergence of our fitting routine near the pole at $b\approx-90$, resulting in posterior distributions ‘pinched’ at the southern pole. We therefore use the rotated Magellanic Stream coordinate frame from Nidever et al. (2008). The $B_{\mathrm{MS}}=0$ latitude of this frame closely follows the LMC orbit shown in Figure 3, and the present-day LMC is located at $L_{\mathrm{MS}}=0^{\circ}$ We have verified that this choice of frame does not influence the best-fit model parameters in any way, but enables more efficient sampling (away from any poles) and well-behaved posterior distributions in our probabilistic model-fitting routine.
@@
We depart from the PP21 model in that we explicitly expand all model parameters in terms of Galactocentric distance, making a tomographic measurement of the reflex motion. We adopt a simple linear expansion of each model parameter, such that the number of parameters is increased by a factor of two. Each parameter can now have different values at $r_{\mathrm{gal}}=40$ and at $r_{\mathrm{gal}}=120$, with a linear trend between those two values. This results in two independent parameter vectors $S_{40}$ and $S_{120}$, each containing 9 parameters: $v_{\mathrm{reflex}}$, $\ell_{\mathrm{MS,apex}}$, $b_{\mathrm{MS,apex}}$, $\langle v_{\mathrm{r}}\rangle$, $\langle v_{\phi}\rangle$, $\langle v_{\theta}\rangle$, ${\sigma_{\mathrm{v,r}}}$, ${\sigma_{\mathrm{v,\ell}}}$, ${\sigma_{\mathrm{v,b}}}$. We only fit stars out to $r_{\mathrm{gal}}=120$ with this model, reducing the sample size to $\approx 800$ stars. Restricting the fit to $r_{\mathrm{gal}}<120$ ensures relatively even coverage across the sky at the upper distance range, since almost all stars beyond this distance are from our southern MagE survey.
For each star, the velocity model is calculated using the model parameters at that distance — linearly evaluated using a given parameter set $\{S_{40},S_{120}\}$ — and the likelihood is computed. This continuity model avoids the pitfalls of binning the data in distance — the substantial distance uncertainties are bound to place stars in incorrect bins, smearing out the signal (see the discussion in YPP24). The continuity model utilizes statistical information from the entire sample at once to uncover the distance-dependence of the model parameters, under the justified assumption that the true parameters vary smoothly and continuously over Galactocentric distance (Garavito-Camargo et al., 2019, 2021).
@@
The likelihood utilizes the full proper motion covariance matrix provided by Gaia, and propagates our measured distance uncertainties into the tangential velocity uncertainties (see Equations 5-10 in PP21). For more details on the model formalism and likelihood, we refer to PP21 and YPP24.
The $v_{\mathrm{travel}}$ continuity model has 18 free parameters: 9 describing the model parameters at $r_{\mathrm{gal}}=40$ kpc, and 9 at $r_{\mathrm{gal}}=120$ kpc. We adopt uninformative priors on all parameters. For the $B_{\mathrm{MS,apex}}$ latitude, a uniform prior on $\cos{[B_{\mathrm{MS,apex}}]}$ is used, and for the intrinsic dispersion parameters a uniform prior on the logarithm of the dispersion is used. An initial maximum-likelihood fit is run using scipy.optimize (Virtanen et al., 2020), and outliers further than $3\sigma$ from the model (accounting for the intrinsic dispersion) are clipped from the sample. Five such iterations are performed, resulting in $\approx 20$ outliers being removed from the sample. These outliers could belong to undetected substructures, or might simply have poorly-fit spectra with spurious parameters. Using the outlier-clipped dataset, the posterior distribution of model parameters is sampled with the nested sampling routine dynesty (Speagle, 2020). The full corner plot of posterior distributions is shown in Appendix B, and the best-fit parameters are summarized in Table 1.
@@
$v_{\mathrm{travel}}$ [km s-1] $5_{-4}^{+5}$ $52_{-12}^{+12}$
$L_{\mathrm{MS,apex}}$ [deg] $-59_{-19}^{+20}$ $-23_{-15}^{+15}$
$B_{\mathrm{MS,apex}}$ [deg] $25_{-18}^{+19}$ $-2_{-14}^{+13}$
$\langle v_{\mathrm{r}}\rangle$ [km s-1] $-11_{-5}^{+5}$ $-23_{-10}^{+9}$
$\langle v_{\mathrm{\phi}}\rangle$ [km s-1] $12_{-4}^{+4}$ $-10_{-8}^{+8}$
$\langle v_{\mathrm{\theta}}\rangle$ [km s-1] $0_{-4}^{+5}$ $6_{-12}^{+12}$
${\sigma_{\mathrm{v,r}}}$ [km s-1] $105_{-4}^{+4}$ $67_{-5}^{+6}$
${\sigma_{\mathrm{v,\ell}}}$ [km s-1] $69_{-3}^{+4}$ $32_{-5}^{+6}$
${\sigma_{\mathrm{v,b}}}$ [km s-1] $54_{-3}^{+3}$ $43_{-6}^{+7}$
@@
To ensure that our model implementation is consistent with previous measurements from PP21, we fit their dataset using our model in Appendix C, reproducing their results. As an additional check to verify our results, we fit all stars in our dataset with a distance-independent 9-parameter $v_{\mathrm{travel}}$ model, directly analogous to that used by PP21 and YPP24. These results are presented in Appendix D, and are broadly consistent with the 18-parameter continuity model results presented in the rest of this section.
@@ 4.2 Reflex Response Results
Figure 6 summarizes the key results from the continuity model measurement of $v_{\mathrm{travel}}$. The top panel illustrates the posterior distribution of the $v_{\mathrm{travel}}$ direction in Galactic coordinates, evaluated at $r_{\mathrm{gal}}=50$ kpc and $r_{\mathrm{gal}}=100$ kpc. Our measured travel velocity relative to the Galactic outskirts points directly at a past location along the LMC’s orbit, close to the present-day location of the LMC. At $r_{\mathrm{gal}}\approx 50$ kpc, the travel velocity is less constrained and points towards an older point in the LMC’s past orbit, just barely in the opposing Galactic quadrant to the LMC. However at larger distances, the travel velocity firmly points closer towards the present-day location of the LMC.
The bottom panel of Figure 6 shows the posterior distribution of the measured amplitude of $v_{\mathrm{travel}}$ as a function of Galactocentric distance $r_{\mathrm{gal}}$, with the $1\sigma$ credible interval shaded. The $v_{\mathrm{travel}}$ amplitude significantly grows as a function of distance, from $\approx 5$ km s-1 at $r_{\mathrm{gal}}=40$ kpc to a maximum of $v_{\mathrm{travel}}=50\pm 20$ at $r_{\mathrm{gal}}=120$ kpc. Along with the recent measurements of YPP24, this is the first measurement of the growing reflex velocity as a function of distance. Whereas YPP24 have a single bin containing all their stars beyond 50 kpc, our larger dataset and continuity model allow us to model the $v_{\mathrm{travel}}$ variation out to $120$ kpc.
The MW+LMC simulations from GC19 are also shown in the bottom panel of Figure 6. Since we only require the maximum-likelihood parameters (and not their uncertainties) from the simulations, we fit the simulations with scipy.optimize (Virtanen et al., 2020) instead of dynesty. We have verified that measuring $v_{\mathrm{travel}}$ directly from the simulation particles — by averaging particle velocities in spherical shells, as done by YPP24 — produces a similar slope, and the linear model is sufficient to capture the overall growth of $v_{\mathrm{travel}}$. However, fitting the simulation with the full model is important to disentangle the mean halo velocity terms from the $v_{\mathrm{travel}}$ amplitude. The bottom panel of Figure 7 demonstrates that the amplitude of the travel velocity $v_{\mathrm{travel}}$ is most consistent with simulations of a $\gtrsim 1.8\times 10^{11}\,M_{\odot}$ LMC from GC19.
@@
The top panel of Figure 7 shows that the GC19 simulations predict a reflex velocity pointing towards a relatively recent position of the LMC’s orbit, consistent with our measurement. On the other hand, the measurements of PP21 and YPP24 point towards the opposing Galactic quadrant. The difference between those measurements and the GC19 simulations is acknowledged in those works (see also Vasiliev 2023). The measurements in YPP24 get closer to the present-day location of the LMC at larger distances, and their appendices explore how different selection effects can non-trivially affect the inferred apex direction. Furthermore, the predicted $v_{\mathrm{travel}}$ amplitude is quite small ($\lesssim 20$ km s-1) for the distances covered by the majority of the PP21 and YPP24 samples, making it more challenging to measure than at larger distances.
@@
Figure 8 offers an alternative perspective of our measurements by showing the Galactocentric components of the $v_{\mathrm{travel}}$ velocity. As a reminder, $v_{\mathrm{travel}}$ is the inferred velocity of the inner galaxy relative to the outer halo, which sources the apparent reflex velocity of those outer halo stars as measured by us. The corresponding $v_{\mathrm{travel}}$ values for each MW+LMC simulation from Garavito-Camargo et al. (2019) are also evaluated at the reference distance of $r_{\mathrm{gal}}=100$ kpc, along with past measurements of $v_{\mathrm{travel}}$ from Vasiliev et al. (2021), PP21, and YPP24. Note that Vasiliev et al. (2021) measure $v_{\mathrm{travel}}$ by fitting the orbit of the Sagittarius stream, whereas PP21 and YPP24 directly fit the kinematics of halo stars.
Figure 8 demonstrates that although all measurements and simulations broadly agree regarding the $v_{\mathrm{Z}}$ velocity (unsurprisingly, since the LMC remains in the southern hemisphere throughout its recent orbit), there is a difference in the $v_{\mathrm{Y}}$ velocity. Specifically, we measure $v_{\mathrm{Y}}\lesssim 0$ in this work (as do Vasiliev et al. 2021), whereas PP21 and YPP24 measure $v_{\mathrm{Y}}>0$. Regardless, this difference is not particularly dramatic, especially relative to the uncertainties in both the past and present measurements.
@@ 4.3 The Global Velocities of the Outer Halo
Another interesting signature in the top panel of Figure 10 is the non-zero azimuthal rotation measured by the $\langle v_{\phi}\rangle$ component. Whereas PP21 and YPP24 measured a net negative $\langle v_{\phi}\rangle$ at all distances, our best-fit model has net positive $\langle v_{\phi}\rangle$ out to $\approx 80$ kpc, after which it switches to being negative. Our relative trend matches YPP24, with $\langle v_{\phi}\rangle$ becoming more negative as a function of distance. In our Galactocentric convention, this corresponds to a net retrograde halo that switches to prograde in the outskirts, although the uncertainties are large, and this measurement is highly sensitive to the assumed Solar motion (see Deason et al. 2017).
Figure 11: Velocity anisotropy parameter $\beta$ as a function of Galactocentric distance, as derived from the velocities shown in Figure 10. Previous measurements are shown for comparison: Cunningham et al. (2019) from the HALO7D survey, Bird et al. (2021) using SDSS/LAMOST K giants, and Han et al. (in prep) using the H3 Survey. Our sample constrains $\beta$ twice as far as previous measurements, suggesting a radially anisotropic halo out to $\gtrsim 100$ kpc.
The bottom panel of Figure 10 shows the radial trends of velocity dispersion fitted alongside our $v_{\mathrm{travel}}$ model. Note that the continuity model assumed a linear trend of the log-dispersion as a function of $r_{\mathrm{gal}}$, producing the apparent curvature of the lines in the bottom panel of Figure 10. We emphasize that these are the intrinsic dispersions from our $v_{\mathrm{travel}}$ model, and are therefore inherently de-convolved from the measurement uncertainties in radial velocities and proper motions. All components of the halo velocity dispersion fall as a function of distance, as has been seen in prior studies mapping the outer halo (e.g., Bird et al., 2021). It is apparent that the radial dispersion is significantly higher than the tangential dispersions across all Galactocentric radii.
@@
This is a crucial parameter for dynamical mass estimates of the MW that use the Jeans formalism (e.g. Binney & Tremaine, 2008; Wang et al., 2020). We compute $\beta$ from $40-120$ kpc using the bulk halo velocities (and their intrinsic dispersions) shown in Figure 10. For example, $\langle v_{r}^{2}\rangle=\langle v_{r}\rangle^{2}+\sigma_{\mathrm{v,r}}^{2}$. Since our intrinsic dispersions are computed in the observed rather than Galactocentric frame, we make the simplifying assumption that $\sigma_{\mathrm{v,r}}\approx\sigma_{\mathrm{v,los}}$, $\sigma_{\mathrm{v,\phi}}\approx\sigma_{\mathrm{v,\ell}}$, and $\sigma_{\mathrm{v,\theta}}\approx\sigma_{\mathrm{v,b}}$ which is suitable for the distant stars in our sample. Our measured Galactocentric profile of $\beta$ is shown in Figure 11, with some corresponding measurements from the literature displayed for comparison.
Simulations of hierarchical galaxy formation predict that at large distances, the stellar halo should have radially-biased orbits with $\beta>0$ (e.g., Diemand et al., 2005; Abadi et al., 2006; Sales et al., 2007; Debattista et al., 2008; Amorisco, 2017; Loebman et al., 2018). We measure a significant radial bias in the motions of stars out to $r_{\mathrm{gal}}\approx 120$ kpc, with a near-constant $\beta\approx 0.6$. These are the most distant all-sky measurements of the halo’s velocity ellipsoid to date, supporting the idea that bulk of the MW halo was hierarchically assembled even in its most distant outskirts. Importantly, along with the reflex motion described earlier, these $\beta$ values should be incorporated in future studies that measure the dynamical mass of the MW using the Jeans equations (see Wang et al., 2020, for a recent review).

### v2

@@ 4.1 Global Velocity Model with Reflex Motion
The direction of $v_{\mathrm{travel}}$ can be encoded by any on-sky angular quantity. Although PP21 used the Galactic coordinates $\ell{}$ and $b$ directly, we find that this leads to inefficient convergence of our fitting routine near the pole at $b\approx-90^{\circ}$, resulting in posterior distributions ‘pinched’ at the southern pole. We therefore use the rotated Magellanic Stream coordinate frame from Nidever et al. (2008). The $B_{\mathrm{MS}}=0^{\circ}$ latitude of this frame closely follows the LMC orbit shown in Figure 3, and the present-day LMC is located at $L_{\mathrm{MS}}=0^{\circ}$. We have verified that this choice of frame does not influence the best-fit model parameters in any way, but enables more efficient sampling (away from any poles) and well-behaved posterior distributions in our probabilistic model-fitting routine.
@@
We depart from the PP21 model in that we explicitly expand all model parameters in terms of Galactocentric distance, making a tomographic measurement of the reflex motion. We adopt a simple linear expansion of each model parameter, such that the number of parameters is increased by a factor of two. Each parameter can now have different values at $r_{\mathrm{gal}}=40$ and at $r_{\mathrm{gal}}=120$, with a linear trend between those two values. This results in two independent parameter vectors $S_{40}$ and $S_{120}$, each containing 9 parameters: $v_{\mathrm{reflex}}$, $\ell_{\mathrm{MS,apex}}$, $b_{\mathrm{MS,apex}}$, $\langle v_{\mathrm{r}}\rangle$, $\langle v_{\phi}\rangle$, $\langle v_{\theta}\rangle$, ${\sigma_{\mathrm{v,r}}}$, ${\sigma_{\mathrm{v,\ell}}}$, ${\sigma_{\mathrm{v,b}}}$. We only fit stars out to $r_{\mathrm{gal}}=120$ with this model, reducing the sample size to $821$ stars — $373$ from H3, $261$ from SEGUE, and $187$ from MagE. Restricting the fit to $r_{\mathrm{gal}}<120$ ensures relatively even coverage across the sky at the upper distance range, since almost all stars beyond this distance are from our southern MagE survey. We have tested three other choices of the fitted distance range ($40-100$ kpc, $30-120$ kpc, $30-100$ kpc), and find that they do not alter the main conclusions of this work. In the $30-120$ kpc fit, the overall amplitude of the fitted $v_{\mathrm{travel}}$ is lower at all distances. We speculate that this is due to the linearity assumption breaking down at $r_{\mathrm{gal}}<40$ kpc, since the $v_{\mathrm{travel}}$ signature is expected to flatten at these distances (GC19, YPP24). The $40-120$ kpc sample has the most homogeneous sky coverage across all distances, and is adopted as our fiducial sample for the rest of this work.
To prevent outliers from biasing our model fit, we clip velocity outliers from our sample. Specifically, we split the sample into 3 distance bins bounded by the following edges: $[40,60,90,120]$ kpc. In each distance bin, we compute the sigma-clipped mean and standard deviation of each velocity component: one radial and two tangential. We remove stars that are $\geq 3.5\sigma$ away from the mean in either velocity component. This results in 34 stars being removed from the sample, the vast majority of which are within $r_{\mathrm{gal}}<60$ kpc. These stars could belong to undetected substructures (or known structures like Sagittarius, missed by our angular momentum selection for Sgr members), or could simply have spurious stellar parameters or velocity measurements. Our fitted sample of stars from $40-120$ kpc therefore has 772 stars.
For each star, the velocity model is calculated using the model parameters at that distance — linearly evaluated using a given parameter set $\{S_{40},S_{120}\}$ — and the likelihood is computed. This continuity model avoids the pitfalls of binning the data in distance. The substantial distance uncertainties are bound to place stars in incorrect bins, smearing out the signal by placing (intrinsically more numerous) nearby stars at larger distances, damping the inferred reflex velocity. The continuity model utilizes statistical information from the entire sample at once to uncover the distance-dependence of the model parameters, under the assumption that the true parameters vary linearly and continuously over Galactocentric distance.
Conceptually, the continuity model reduces the variance at any given distance, while potentially increasing the bias — for example, the parameters at nearby distances are by definition influenced by data at larger distances. By applying our model to the simulations from GC19, we have verified that the linear continuity model is sufficient to capture the overall trend of the $v_{\mathrm{travel}}$ parameters, as well as the velocity dispersion parameters. In Appendix C we fit a simulation that matches the size and spatial sampling of our dataset, affirming that our methodology is robust against uneven spatial sampling.
@@
The likelihood utilizes the full proper motion covariance matrix provided by Gaia, and propagates our measured distance uncertainties into the tangential velocity uncertainties (see Equations 5-10 in PP21). The uncertainties of the on-sky positions are assumed to be negligible. For more details on the model formalism and likelihood, we refer to PP21 and YPP24.
The $v_{\mathrm{travel}}$ continuity model has 18 free parameters: 9 describing the model parameters at $r_{\mathrm{gal}}=40$ kpc, and 9 at $r_{\mathrm{gal}}=120$ kpc. We adopt uninformative priors on all parameters. For the $B_{\mathrm{MS,apex}}$ latitude, a uniform prior on $\cos{[B_{\mathrm{MS,apex}}]}$ is used, and for the intrinsic dispersion parameters a uniform prior on the logarithm of the dispersion is used. An initial maximum-likelihood fit is run using scipy.optimize (Virtanen et al., 2020). The posterior distribution of model parameters is sampled with the nested sampling routine dynesty (Speagle, 2020). The full corner plot of posterior distributions is shown in Appendix F, and the best-fit parameters are summarized in Table 1. The best-fit parameters from the linear continuity model are also evaluated at a reference distance of $r_{\mathrm{gal}}=100$ kpc, and listed in Table 2.
@@
$v_{\mathrm{travel}}$ [km s-1] $6_{-4}^{+5}$ $47_{-12}^{+12}$
$\ell_{\mathrm{MS,apex}}$ [deg] $-77_{-22}^{+21}$ $-10_{-15}^{+17}$
$b_{\mathrm{MS,apex}}$ [deg] $12_{-19}^{+20}$ $3_{-14}^{+14}$
$\langle v_{\mathrm{r}}\rangle$ [km s-1] $-8_{-5}^{+5}$ $-25_{-10}^{+9}$
$\langle v_{\mathrm{\phi}}\rangle$ [km s-1] $14_{-5}^{+5}$ $-14_{-8}^{+8}$
$\langle v_{\mathrm{\theta}}\rangle$ [km s-1] $-1_{-5}^{+5}$ $-3_{-12}^{+12}$
${\sigma_{\mathrm{v,r}}}$ [km s-1] $106_{-4}^{+4}$ $69_{-5}^{+6}$
${\sigma_{\mathrm{v,\ell}}}$ [km s-1] $78_{-4}^{+4}$ $20_{-4}^{+5}$
${\sigma_{\mathrm{v,b}}}$ [km s-1] $56_{-3}^{+3}$ $33_{-5}^{+6}$
@@
To ensure that our model implementation is consistent with previous measurements from PP21, we fit their dataset using our model in Appendix A, reproducing their results. As an additional check to verify our results, we fit all stars in our dataset with a distance-independent 9-parameter $v_{\mathrm{travel}}$ model, directly analogous to that used by PP21 and YPP24. These results are presented in Appendix E, and are broadly consistent with the 18-parameter continuity model results presented in the rest of this section.
Table 2: Summary of the best-fit parameters from our $v_{\mathrm{travel}}$ continuity model (see Table 1), linearly evaluated at a reference distance of $100$ kpc. For convenience and ease of comparison, we also provide the apex direction transformed to Galactic coordinates, and the $v_{\mathrm{travel}}$ velocity in Galactocentric coordinates (see Figure 9).
         Parameter          $r_{\mathrm{gal}}=100$ kpc
$v_{\mathrm{travel}}$ [km s-1] $37_{-9}^{+9}$
$\ell_{\mathrm{MS,apex}}$ [deg] $-27_{-10}^{+11}$
$b_{\mathrm{MS,apex}}$ [deg] $6_{-9}^{+9}$
$\langle v_{\mathrm{r}}\rangle$ [km s-1] $-21_{-7}^{+6}$
$\langle v_{\mathrm{\phi}}\rangle$ [km s-1] $-7_{-6}^{+6}$
$\langle v_{\mathrm{\theta}}\rangle$ [km s-1] $-2_{-9}^{+9}$
${\sigma_{\mathrm{v,r}}}$ [km s-1] $77_{-4}^{+5}$
${\sigma_{\mathrm{v,\ell}}}$ [km s-1] $28_{-4}^{+5}$
${\sigma_{\mathrm{v,b}}}$ [km s-1] $38_{-4}^{+4}$
$\ell_{\mathrm{apex}}$ [deg] $-80_{-18}^{+18}$
$b_{\mathrm{apex}}$ [deg] $-59_{-9}^{+11}$
$v_{\mathrm{travel,x}}$ [km s-1] $3_{-5}^{+5}$
$v_{\mathrm{travel,y}}$ [km s-1] $-17_{-5}^{+6}$
$v_{\mathrm{travel,z}}$ [km s-1] $-31_{-9}^{+9}$
@@ 4.2 Reflex Response Results
Figure 6 summarizes the key results from the continuity model measurement of $v_{\mathrm{travel}}$. The top panel illustrates the posterior distribution of the $v_{\mathrm{travel}}$ direction in Galactic coordinates, evaluated at the reference distance of $r_{\mathrm{gal}}=100$ kpc. As a reminder, the continuity model evaluated at this reference distance utilizes information from stars spanning $r_{\mathrm{gal}}=40-120$ kpc. Our measured travel velocity relative to the Galactic outskirts points directly at a past location along the LMC’s orbit, close to the present-day location of the LMC.
The bottom panel of Figure 6 shows the posterior distribution of the measured amplitude of $v_{\mathrm{travel}}$ as a function of Galactocentric distance $r_{\mathrm{gal}}$, with the $1\sigma$ credible interval shaded. The $v_{\mathrm{travel}}$ amplitude significantly grows as a function of distance, from $\approx 6$ km s-1 at $r_{\mathrm{gal}}=40$ kpc to a maximum of $v_{\mathrm{travel}}=47\pm 12$ at $r_{\mathrm{gal}}=120$ kpc. Along with the recent measurements of YPP24, this is the first measurement of the growing reflex velocity as a function of distance. Whereas YPP24 have a single bin containing all their stars beyond 50 kpc, our larger dataset and continuity model allow us to model the $v_{\mathrm{travel}}$ variation out to $120$ kpc.
The MW+LMC simulations from GC19 are also shown in the bottom panel of Figure 6. Since we only require the maximum-likelihood parameters (and not their uncertainties) from the simulations, we fit the simulations with scipy.optimize (Virtanen et al., 2020) instead of dynesty. We have verified that measuring $v_{\mathrm{travel}}$ directly from the simulation particles — by averaging particle velocities in spherical shells, as done by YPP24 — produces a similar slope, and the linear model is sufficient to capture the overall growth of $v_{\mathrm{travel}}$. However, fitting the simulation with the full model is important to disentangle the mean halo velocity terms from the $v_{\mathrm{travel}}$ amplitude. The bottom panel of Figure 8 demonstrates that the amplitude of the travel velocity $v_{\mathrm{travel}}$ is most consistent with simulations of a $\gtrsim 1.8\times 10^{11}\,M_{\odot}$ LMC from GC19, or an LMC that is $15\%$ the mass of the Milky Way.
Figure 7: The apex direction of $v_{\mathrm{travel}}$ measured in this work (red lines) as a function of Galactocentric radius $r_{\mathrm{gal}}$ with $1\sigma$ uncertainties shaded. The dashed blue line shows the same quantities from the $1.8\times 10^{11}\,M_{\odot}$ LMC3 simulation of GC19.
Figure 7 shows the apex direction of $v_{\mathrm{travel}}$ as a function of Galactocentric radius. The $1.8\times 10^{11}\,M_{\odot}$ LMC simulation from GC19 is shown for comparison. In both the data and simulations, there is little evolution in the apex direction as a function of distance.
@@
The top panel of Figure 8 shows that the GC19 simulations predict a reflex velocity pointing towards a relatively recent position of the LMC’s orbit, consistent with our measurement. On the other hand, the measurements of PP21 and YPP24 point towards the opposing Galactic quadrant. The difference between those measurements and the GC19 simulations is acknowledged in those works (see also Vasiliev 2023). The measurements in YPP24 get closer to the present-day location of the LMC at larger distances, and their appendices explore how different selection effects can non-trivially affect the inferred apex direction. Furthermore, the predicted $v_{\mathrm{travel}}$ amplitude is quite small ($\lesssim 20$ km s-1) for the distances covered by the majority of the PP21 and YPP24 samples, making it more challenging to disentangle from other bulk motions in the halo than at larger distances.
@@
Figure 9 offers an alternative perspective of our measurements by showing the Galactocentric components of the $v_{\mathrm{travel}}$ velocity. As a reminder, $v_{\mathrm{travel}}$ is the inferred velocity of the inner galaxy relative to the outer halo, which sources the apparent reflex velocity of those outer halo stars as measured by us. The corresponding $v_{\mathrm{travel}}$ values for each MW+LMC simulation from Garavito-Camargo et al. (2019) are also evaluated at the reference distance of $r_{\mathrm{gal}}=100$ kpc, along with past measurements of $v_{\mathrm{travel}}$ from Vasiliev et al. (2021), PP21, and YPP24. Note that Vasiliev et al. (2021) measure $v_{\mathrm{travel}}$ by fitting the orbit of the Sagittarius stream, whereas PP21 and YPP24 directly fit the kinematics of halo stars. The Vasiliev et al. (2021) measurement in Figure 9 was provided by E. Vasiliev (private communication).
Figure 9 demonstrates that although all measurements and simulations broadly agree regarding the $v_{\mathrm{Z}}$ velocity (unsurprisingly, since the LMC remains in the southern hemisphere throughout its recent orbit), there is a difference in the $v_{\mathrm{Y}}$ velocity. Specifically, we measure $v_{\mathrm{Y}}\lesssim 0$ in this work (as do Vasiliev et al. 2021), whereas PP21 and YPP24 measure $v_{\mathrm{Y}}>0$. This difference is modest relative to the uncertainties in both the past and present measurements. In Appendix A, we investigate this discrepancy in detail, finding that the measurements in PP21 and YPP24 were likely influenced by the presence of Sagittarius Stream stars in their datasets (see also Byström et al. 2024). We conclude that the difference between our measurement and past measurements from PP21 and YPP24 is primarily due to the different selection cuts used to remove Sagittarius stars from the datasets.
@@ 4.3 The Global Velocities of the Outer Halo
Another interesting signature in the top panel of Figure 11 is the non-zero azimuthal rotation measured by the $\langle v_{\phi}\rangle$ component. Whereas PP21 and YPP24 measured a net negative $\langle v_{\phi}\rangle$ at all distances, our best-fit model has net positive $\langle v_{\phi}\rangle$ out to $\approx 80$ kpc, after which it switches to being negative. Our relative trend matches YPP24, with $\langle v_{\phi}\rangle$ becoming more negative as a function of distance. In our Galactocentric convention, this corresponds to a net retrograde halo that switches to prograde in the outskirts, although the uncertainties are large (see Deason et al. 2017).
The bottom panel of Figure 11 shows the radial trends of velocity dispersion fitted alongside our $v_{\mathrm{travel}}$ model. Note that the continuity model assumed a linear trend of the log-dispersion as a function of $r_{\mathrm{gal}}$, producing the apparent curvature of the lines in the bottom panel of Figure 11. We emphasize that these are the intrinsic dispersions from our $v_{\mathrm{travel}}$ model, and are therefore inherently de-convolved from the measurement uncertainties in radial velocities and proper motions. All components of the halo velocity dispersion fall as a function of distance, as has been seen in prior studies mapping the outer halo (e.g., Bird et al., 2021). It is apparent that the radial dispersion is significantly higher than the tangential dispersions across all Galactocentric radii. We have verified that directly calculating the velocity dispersion in radial bins — and subtracting the median measurement error in each bin, in quadrature — produces quantitatively similar profiles. We do note that it is challenging to measure the intrinsic tangential dispersions at large distances, since they are sensitive to the adopted proper motion and distance measurement uncertainties.
@@
This is a crucial parameter for dynamical mass estimates of the MW that use the Jeans formalism (e.g. Binney & Tremaine, 2008; Wang et al., 2020). We compute $\beta$ from $40-120$ kpc using the bulk halo velocities (and their intrinsic dispersions) shown in Figure 11. For example, $\langle v_{r}^{2}\rangle=\langle v_{r}\rangle^{2}+\sigma_{\mathrm{v,r}}^{2}$. Since our intrinsic dispersions are computed in the observed rather than Galactocentric frame, we make the simplifying assumption that $\sigma_{\mathrm{v,r}}\approx\sigma_{\mathrm{v,los}}$, $\sigma_{\mathrm{v,\phi}}\approx\sigma_{\mathrm{v,\ell}}$, and $\sigma_{\mathrm{v,\theta}}\approx\sigma_{\mathrm{v,b}}$ which is suitable for the distant stars in our sample. We have verified by fitting the simulations of Garavito-Camargo et al. (2019) that we can recover their input $\beta(r_{\mathrm{gal}})$ profile using our fitting methodology. Our measured Galactocentric profile of $\beta$ is shown in Figure 12, with some corresponding measurements from the literature displayed for comparison.
Figure 12: Velocity anisotropy parameter $\beta$ as a function of Galactocentric distance, as derived from the velocities shown in Figure 11. Previous measurements are shown for comparison: Cunningham et al. (2019) from the HALO7D survey, Bird et al. (2021) using SDSS/LAMOST K giants, and Han et al. (2024) using the H3 Survey. Our sample constrains $\beta$ twice as far as previous measurements, suggesting a radially anisotropic halo out to $\gtrsim 100$ kpc.
Simulations of hierarchical galaxy formation predict that at large distances, the stellar halo should have radially-biased orbits with $\beta>0$ (e.g., Diemand et al., 2005; Abadi et al., 2006; Sales et al., 2007; Debattista et al., 2008; Amorisco, 2017; Loebman et al., 2018). We measure a significant radial bias in the motions of stars out to $r_{\mathrm{gal}}\approx 120$ kpc, with $\beta$ ranging from $\approx 0.6-0.8$. These are the most distant all-sky measurements of the halo’s velocity ellipsoid to date, supporting the idea that bulk of the MW halo was hierarchically assembled even in its most distant outskirts. Importantly, along with the reflex motion described earlier, these $\beta$ values should be incorporated in future studies that measure the dynamical mass of the MW using the Jeans equations (see Wang et al., 2020, for a recent review).